Express 3+4i in polar form and hence determine the exact square roots of 3+4i in
the form a + ib, where a, b are real numbers.

Answer by stanbon(75887)

Express 3+4i in polar form and hence determine the exact square roots of 3+4i in
the form a + ib, where a, b are real numbers.
--------
r = sqrt[3^2+4^2] = sqrt[25] = 5
----
angle = arctan(4/3) = 36.87 degrees
-----------------------
So, 3+4i = 5[cis(36+87+360n)]
-----
sqrt(3+4i) = ?
Let n = 0, then sqrt(3+4i) = sqrt(5)[cis(36.87/2)]
Let n = 1, then sqrt(3+4i) = sqrt(5)[cis[(36.87+360)/2]
-----------------
Cheers,
Stan H.
===================